Ticket Number: HEA-NOC/20090209-3 Ticket Status: CLOSE
Ticket Type: unscheduled Resolver: HEA-NOC
Ticket Opened: 20090209 11:28 UTC Problem Start: 20090209 08:49 UTC
Ticket Closed: 20090211 14:50 UTC Problem End: 20090209 18:21 UTC
Site/line: Mater Dei
Several flaps to Mater Dei and St Pat's from 08:49 to 09:10
Loss of connectivity to Mater Dei and SPD.
20090209 11:15 BMA Rang TCD PoP, Ken reports some flooding and will
20090209 13:29 BMA Paul J rang from TCD - no issues on their side. Likely
source of problem at Mater Dei, but their phone systems appear to be down
at present. Asked TCD to contact them.
20090209 15:22 BMA DCU confirm power issues at Mater Dei, whom SPD are
connected through. Closing ticket.
20090209 17:07 BMA Further outage at 17:01. DCU indicate it may be
tomorrow before power is stable again. Leaving ticket open to monitor.
20090211 14:45 BMA ESB fixed power at Mater Dei. Switch came back up at
18:21 on 20090209, no outages since. GE1/0/1 is still suffering from
recognised problem of low light on the extended DCC ring. Closing ticket.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20090209-3
HEAnet Limited. Registered in Ireland, No. 275301.